# Machine Learning Interatomic Potentials Enable Molecular Dynamics Simulations of Doped MoS2

### v1

@@ III.2 Demonstration of the MLIP
Analysis of the atom trajectories from the N-doped MoS2 simulation shows two behaviors. First, many of N dopant atoms chemically react with the MoS2 to form Mo-S-N complexes, as shown in Figure 10(i). Both substitutional and intercalated dopants exhibit this behavior and, once the complexes form, they are very stable such that the dopant atoms have very low diffusivity. However, some N atoms exhibits long diffusion paths, as shown in Figure 10(ii). All instances of this long-range motion occur in pairs where two N atoms move together as a unit. This coordinated movement indicates the formation of N₂ molecules within the MoS₂ matrix, demonstrating the tendency of nitrogen to maintain its diatomic molecular character even when incorporated as a dopant. This shows that the MLIP simulations are effectively able to capture the chemical bond formation and compound synthesis characteristic of MoS2 with non-metal dopants.
IV Conclusions
This study demonstrates the first application of machine learning interatomic potentials for investigating doped MoS2 systems across 25 elements spanning the periodic table. By leveraging Meta’s UMA, we successfully bridged the accuracy-efficiency gap between density functional theory and molecular dynamics, enabling large-scale simulations.
@@ IV Conclusions
Our benchmarking established the reliability of UMA potentials for doped MoS2 systems, with mean absolute errors of 0.374 eV for UMA-small and 0.404 eV for UMA-medium in formation energy predictions compared to DFT reference calculations. Additionally, through local structure analysis, we showed that the UMA small MLIP is able to capture the lattice distortions caused by the dopants for most cases. This confirmed that the models accurately capture the structural effects of doping across diverse chemical elements. Our tests also highlighted the limitations of the UMA dataset and identified cases where the MLIP could be finetuned for improved accuracy.
@@
To demonstrate the MLIP, we ran simulations of 3,100-atom doped MoS2 systems using UMA small. The heating-cooling molecular dynamics simulations revealed four distinct dopant behaviors in MoS2: clustering metals that aggregated during thermal treatment and could induce layer fracturing, non-clustering metals that maintained mobility without aggregation, light diffusive metals that exhibited through-layer diffusion creating continuous dopant flow channels, and chemically reactive non-metals that form stable molecular compounds within the MoS2 matrix. Analysis of representative examples of these groups showed that the simulation could capture complex phenomena including dopant clustering, interlayer diffusion, chemical compound formation, and structural modifications. The findings provide fundamental insight into dopant-host interactions that govern the performance of doped MoS2 in tribological, electronic, and optoelectronic applications. Future work will extend the computational framework to investigate dopant concentration effects, temperature-dependent phase behavior, and the influence of external forces on doped MoS2 properties.
Generally, this work establishes a transferable computational framework with open-source implementation that enables high-throughput screening of dopant candidates and optimization of compositions for targeted applications. The methodology presented here is readily applicable to other two-dimensional materials systems, opening new avenues for computational materials discovery in the emerging field of atomically engineered nanomaterials.
Author Contributions
AF conceptualized, coded and performed the simulations and analysis, and wrote the manuscript. AM supervised and guided AF through the process and edited the manuscript.
Conflicts of interest
There are no conflicts to declare.
Funding
This work was supported by the U.S. Department of Energy, Basic Energy Science under Award number DE-SC0024239. Computational resources were provided by the National Science Foundation (NSF) through grant no. 2346744 (CENVAL-ARC) and grant no. 2019144 (Pinnacles).
Acknowledgments
@@
We would like to thank Daniel Miliate for his technical assistance and review. Dr. Fabrice Roncoroni and Dr. David Prendergast from LBNL for technical assistance. Dr. N. Scott Bobbitt, Dr. Michael Chandross and Dr. John F Curry from SNL for providing feedback on the work.
Data Availability
The DFT and MLIP data generated in this paper was produced using codes available on GitHub: Machine-Learning-Interatomic-Potential-UMA-Enables-MD-Simulations-of-MoS2-Doped
Supplementary Information
Machine Learning Interatomic Potential (UMA) Enables MD Simulations of MoS₂ Doped Across the Periodic Table for Tribological, Electronic, and Optoelectronic Applications
Abrar Faiyad1 and Ashlie Martini1
1University of California Merced, Merced, California 95343, United States
amartini@ucmerced.edu
S1 MSD Analysis
To quantify atomic mobility and dopant diffusion in MoS2, mean square displacement (MSD) analysis was performed on the NVT ensemble trajectories from the heating cycle at 1000 K. Trajectory data were processed using custom Python scripts based on the Atomic Simulation Environment (ASE).
The MSD for each atomic species, was calculated as
$\mathrm{MSD}(t)=\frac{1}{N}\sum_{i=1}^{N}\left\langle\left|\mathbf{r}_{i}(t)-\mathbf{r}_{i}(0)\right|^{2}\right\rangle$ (S1)
where $N$ is the number of atoms for the considered species, $\mathbf{r}_{i}(t)$ is the instantaneous position of atom $i$ at time $t$, and the brackets denote averaging over all atoms of a given type. Reference positions were taken from the first NVT production frame. MSD curves were computed for all major species (Mo, S, and the dopant element) by averaging over respective atomic indices. To compare dopant mobility, the slope of the dopant MSD curve was extracted by linearly fitting the last 50% of the trajectory using least-squares regression:
$\mathrm{MSD}(t)\approx mt+b$ (S2)
where the fit parameter $m$ provides the effective mobility of the dopant under constant thermodynamic conditions.
For the total system, long-time diffusive behavior was quantified by estimating the diffusion coefficient $D$ from the Einstein relation,
$\mathrm{MSD}(t)=6Dt+\mathrm{offset}$ (S3)
using the slope of a linear fit to the total MSD in the diffusive regime. All scripts for MSD analysis and visualization are provided in the project repository.
Figure S1: MAE based on dopant type (metal or non-metal) for the UMA small forcefield, Metals show lower error than non-metals.
Figure S2: Formation energies calculated from DFT for different dopants.
Figure S3: Formation energies calculated from UMA small for different dopants.
(a) Ag doped system
(b) Al doped system
(c) Au doped system
(d) C doped system
(e) Cl doped system
(f) Cu doped system
(g) F doped system
(h) Fe doped system
Figure S4: MSD analysis (1/4): Ag–Fe, alphabetical.
(a) Ir doped system
(b) Li doped system
(c) N doped system
(d) Na doped system
(e) Nb doped system
(f) O doped system
(g) Pd doped system
(h) Pt doped system
Figure S5: MSD analysis (2/4): Ir–Pt, alphabetical.
(a) Re doped system
(b) Rh doped system
(c) Ru doped system
(d) Si doped system
(e) Ta doped system
(f) Te doped system
(g) Ti doped system
(h) V doped system
Figure S6: MSD analysis (3/4): Re–V, alphabetical.
(a) Zn doped system
Figure S7: MSD analysis (4/4): Zn.
(a) Undoped system after heating-cooling cycle
Figure S8: Undoped system after heating-cooling cycle

### v2

@@ Methods
Analysis of the atom trajectories from the N-doped MoS2 simulation shows two behaviors. First, many of N dopant atoms chemically react with the MoS2 to form Mo-S-N complexes, as shown in Figure 10(i). Both substitutional and intercalated dopants exhibit this behavior and, once the complexes form, they are very stable such that the dopant atoms have very low diffusivity. However, some N atoms exhibits long diffusion paths, as shown in Figure 10(ii). All instances of this long-range motion occur in pairs where two N atoms move together as a unit. This coordinated movement indicates the formation of N2 molecules within the MoS2 matrix, demonstrating the tendency of nitrogen to maintain its diatomic molecular character even when incorporated as a dopant. This shows that the MLIP simulations are effectively able to capture the chemical bond formation and compound synthesis characteristic of MoS2 with non-metal dopants.
Conclusions
@@
This study demonstrates the first application of universal machine learning interatomic potentials for doped MoS2 systems spanning the periodic table. By comparing the MLIP-predicted formation energy and local structure with calculations from DFT, we evaluated the accuracy and reliability of the UMA potentials for doped MoS2 systems. Our results show that UMA-small has a mean absolute errors of 0.374 eV and UMA-medium 0.404 eV in formation energy predictions compared to DFT reference calculations. Additionally, through local structure analysis, we showed that the UMA small MLIP is able to capture the lattice distortions caused by the dopants for most cases. This confirmed that the models accurately capture the structural effects of doping across many chemical elements. Our tests also highlighted the limitations of the UMA dataset and identified cases where the MLIP could be finetuned for improved accuracy.
@@
To demonstrate the MLIP, we ran simulations of doped MoS2 systems with 3̃,100-atom using UMA small. The heating-cooling molecular dynamics simulations revealed four distinct dopant behaviors in MoS2: clustering metals that aggregated during thermal treatment and could induce layer fracturing, non-clustering metals that maintained mobility without aggregation, light diffusive metals that exhibited through-layer diffusion creating continuous dopant flow channels, and chemically reactive non-metals that form stable molecular compounds within the MoS2 matrix. Analysis of representative examples from these groups showed that the simulation could capture complex phenomena including dopant clustering, interlayer diffusion, chemical compound formation, and structural modifications. The findings provide fundamental insight into dopant-host interactions that govern the performance of doped MoS2 in tribological, electronic, and optoelectronic applications.
Future work will extend the computational framework to investigate dopant concentration effects, temperature dependent phase behavior, and the influence of external force on doped MoS2, as well as to include other doped TMDs.
Acknowledgements
@@
The authors thank Daniel Miliate for his technical assistance and review. Dr. Fabrice Roncoroni and Dr. David Prendergast from LBNL for technical assistance. Dr. N. Scott Bobbitt, Dr. Michael Chandross and Dr. John F Curry from SNL for providing feedback on the work.
Supporting information
  • Supplymentary_Information.pdf: MSD calculation methods, MAE based on dopant type (metal or non-metal), Formation energy calculated from DFT and MLMD, Individual dopant MSD and undoped system after heating-cooling cycle